Rahman is considering to independent projects, Project A and Project B. The


initial cash outlay associated with Project A is RM50,000 and the project


would generate cash flows of RM12,000 per year for six years. While the


initial outlay for Project B would be RM70,000 with generate cash flows of


RM13,000 per year for ten years. The discount rate on both projects is 12


percent. Calculate Internal Rate of Return (IRR) for each project and indicate


if the project should be accepted or not

1
Expert's answer
2022-03-20T19:09:26-0400

"IRR=\frac{PVCOF}{A}"


Project A

Initial investment for project A is RM50,000


IRR="\frac{50000}{12000}" =4.1667


Check 4.1667 in the annuity table under the 6th year.

It falls between 10%@ 4.3553 and 12%@4.1114


"\frac{IRR-Lower limit}{Upper limit-Lower limit}"


"\frac{IRR-10}{12-10}"="\frac{4.1667-4.3553}{4.1114-4.3553}"


"\frac{IRR-10}{2}" =0.7733


IRR=11.55"\%"


PROJECT B

Initial investment is RM 70,000

Cashflows RM13,000 for 10 years


IRR="\frac{70,000}{13,000}" =5.3846


5.3846 under 10th year from the annuity table;

Falls between14"\%" @5.2161and 12"\%" @5.6502


"\frac{IRR-12}{14-12}"="\frac{5.3846-5.6502}{5.2161-5.65002}"


"\frac{IRR-12}{2}"=0.6118


IRR=13.22"\%"


Decision criteria

Cost of capital of capital is12"\%"

If:

IRR<cost of capital, reject the project

IRR>Cost of capital, accept the project

IRR=Cost of capital, investor is indifference


IRR for project A is 11.55"\%" which is less than the cost of capital and this implies that we should reject the project.

IRR for project B is 13.22"\%" which is greater than the cost of capital. Therefore, we should accept the project.